OpenStudy (dls):

A bird flies for 4 seconds with a velocity of |t-2|m/s in a straight line.It covers a distance of ____m and displacement of _____ m.Please help!

OpenStudy (dls):

@mukushla remember...? :| @Yahoo! @experimentX @Vaidehi09 @rdx

OpenStudy (anonymous):

just intergrate the V

OpenStudy (dls):

tell me what's ur answer first!

OpenStudy (anonymous):

0

OpenStudy (dls):

distance

OpenStudy (anonymous):

displacement =0

OpenStudy (dls):

that is obvious,distance?

OpenStudy (ghazi):

is it 16 m

OpenStudy (dls):

no

OpenStudy (anonymous):

0..i think

OpenStudy (dls):

noway

OpenStudy (experimentx):

find the area under this curve for distance |t-2| the displacement is 2

OpenStudy (dls):

no

OpenStudy (dls):

we don't have to integrate

OpenStudy (anonymous):

distance is 48m

OpenStudy (experimentx):

oh ... sorry ... it's velocity.

OpenStudy (dls):

no @Yahoo!

OpenStudy (anonymous):

it is modulus...t-2

OpenStudy (experimentx):

is it both 4m ??

OpenStudy (dls):

history of this question XD

OpenStudy (anonymous):

so it still remains that velocity = t-2 and speed = |t-2| ....

OpenStudy (experimentx):

i though distance was function of time ...yahoo had asked similar Q ... just integrate that abs(t-2) from 0 to 4

OpenStudy (anonymous):

it is the conclusion from the old history

OpenStudy (anonymous):

but @experimentX but that would give answer as 0

OpenStudy (experimentx):

well ... if your displacement was 2 to begin with ... depends on where you choose to measure it ... but wrt initial (it's zero) ... it were function of distance.

OpenStudy (anonymous):

velocity = t-2 for t>= 2 And -t+2 for t<2

OpenStudy (anonymous):

Yup......

OpenStudy (anonymous):

@sauravshakya gud approach

OpenStudy (anonymous):

The answer should be 8..i think

OpenStudy (anonymous):

Oh.. its 4m

OpenStudy (experimentx):

why don't you draw graphs!! ... this is just area of two triangles.